# A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

### v3

@@ Abstract.
The Rust programming language restricts aliasing to provide static safety guarantees. However, in certain situations, developers need to bypass these guarantees by using a set of unsafe features. If they are used incorrectly, these features can reintroduce the types of safety issues that Rust was designed to prevent. We seek to understand how current development tools can be improved to better assist developers who find it necessary to interact with unsafe code. To that end, we study how developers reason about foreign function calls, the limitations of the tools that they currently use, their motivations for using unsafe code, and how they reason about encapsulating it. We conducted a mixed-methods investigation consisting of semi-structured interviews with 19 developers, followed by a survey that reached an additional 160 developers. Our participants were motivated to use unsafe code when they perceived that there was no alternative, and most avoided using it. However, limited tooling support for foreign function calls made participants uncertain about their design choices, and certain foreign aliasing and concurrency patterns were difficult to encapsulate. To overcome these challenges, Rust developers need verification tools that can provide guarantees of soundness within multi-language applications.
Rust, unsafe, memory safety, threat safety, interoperation, foreign functions, motivations, tooling, interview, survey, mixed-methods
copyright: noneccs: Software and its engineering
@@ 1. Introduction
The Rust programming language has made rapid headway as a safe alternative for systems programming due to its static safety guarantees. However, it is still the case that critical systems are written in languages that do not statically prevent memory and thread safety issues. Developers who interoperate with these programs must use Rust’s unsafe features, which bypass its safety restrictions. To minimize the risks of using these features, the Rust community advocates for keeping unsafe code minimal, well documented, and hidden beneath a safe interface (Astrauskas et al., 2020). However, these practices alone are not enough to ensure that Rust fulfills its promise of static safety. Errors in “safe” encapsulations of unsafe implementations have been a significant source of bugs and security vulnerabilities in the Rust ecosystem (Qin et al., 2020; Xu et al., 2021; Zheng et al., 2023). Applications that rely on interoperation or direct access to system resources may expose developers to sources of unsafety that cannot easily be minimized. Making matters worse, Rust’s semantics are still evolving, so few tools are capable of fully testing or verifying these applications.
@@
Interoperation is a common use case for unsafe code (Evans et al., 2020; Astrauskas et al., 2020; Fulton et al., 2021; Höltervennhoff et al., 2023), but most Rust-specific development tools lack support for foreign function calls. Safe interoperation will be crucial to Rust’s future, since both industry and regulatory stakeholders have advocated for adopting Rust as a replacement for existing systems languages (Wallach, 2024; Rebert and Kern, 2024). Automated translation tools are increasingly effective (Larsen, 2018), but for the moment, interoperation is the most practical solution for adopting Rust within preexisting, large-scale systems (Rebert and Kern, 2024). It can be challenging to use Rust correctly in this context, because it has significant differences from other languages (Höltervennhoff et al., 2023; McCormack et al., 2024). A detailed account of these challenges is missing, and we need this to be able to develop effective tooling for safe interoperation. To provide this perspective, we ask the following research questions:
@@
RQ1 (Interoperation). How do Rust developers reason about memory safety across foreign function boundaries? RQ2 (Tooling). What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved?
@@
To design effective interventions for developers who use unsafe code, we need to understand what motivates them to use these features and how they reason about the soundness of their designs. However, most large-scale studies of these topics have analyzed patterns in source code instead of directly engaging with developers (Evans et al., 2020; Astrauskas et al., 2020). Interviews and surveys have provided a missing qualitative perspective, but they tend to have a limited focus on unsafe code, a small sample size, or different goals (Fulton et al., 2021; Höltervennhoff et al., 2023). None of these studies have directly investigated how developers use foreign functions—we aim to fill this knowledge gap. In addition, to serve as guidance for the design of future tools, we seek to generalize existing qualitative results to a broader population of developers. To that end, we ask the following additional research questions:
RQ3 (Motivations). What are Rust developers’ motivations for using unsafe code? RQ4 (Encapsulation). How do Rust developers reason about encapsulating unsafe code?
@@
We provide the results of an exploratory, mixed method study that addresses the research questions above. We conducted semi-structured interviews with 19 developers and used the results to create a community survey that had 160 valid responses. The majority of our participants used unsafe code to call foreign functions, and they described several aliasing and concurrency patterns that made interoperation difficult. Many developers had used Miri, an interpreter, to find bugs, but most were deterred from using it again due to its slow performance and lack of support for key features. Participants prioritized keeping unsafe code minimal, encapsulated, and well documented, but they rarely audited their dependencies and relied on ad-hoc reasoning to justify their decisions. This correctness gap can be addressed by creating verification and bug finding tools that have robust support for multi-language applications and are capable of providing guarantees of soundness.
@@
In Section 0, we describe the role of unsafe code in the Rust ecosystem and describe the categories of tools that Rust developers can use for testing, bug finding, and verification. In Section 0, we provide an overview of related studies of unsafe Rust. We describe our methodology in Section 0. We present our results in Section 0 and discuss their implications in Section 0. We indicate threats to validity in Section 0 before concluding in Section 0. Our replication package is available via GitHub [^1] https://github.com/icmccorm/unsafe-study and FigShare. [^2] https://figshare.com/s/aacea766be06aa6deb97
@@ 2. Background
Rust’s safety guarantees build on the notion of ownership. Values with “copy” semantics do not have an owner and can be copied freely, while values with “move” semantics have a unique owner. Assignment transfers ownership from one alias to another, but ownership can also be borrowed by creating a reference. Each reference has a lifetime, which is the duration of the program where it is used. Rust’s borrow checker uses this property to enforce that a value can have either a unique, mutable reference or many immutable references, but not both simultaneously (Crichton, 2020). A value’s type can implement traits, which define how it behaves. Certain traits affect a value’s capability to be owned and borrowed. The Copy trait gives a value copy semantics, Send allows values to be transferred between threads, and Sync allows multiple threads to reference a value.
@@
Rust’s aliasing rules are conservative and are not universally compatible with certain idioms of systems programming. When these restrictions become burdensome, developers can use the unsafe keyword to enable features that escape the constraints of Rust’s borrow checker. Within an unsafe context, users can dereference raw pointers, execute inline assembly instructions, modify static mutable state, implement unsafe traits, and call unsafe functions (Klabnik and Nichols, 2022). There are a variety of use cases for these features (Astrauskas et al., 2020). In certain situations, unsafe code can be necessary to implement design patterns that are fundamentally incompatible with the restrictions of safe references, such as shared mutable state. Foreign function calls, inline assembly, and intrinsics allow developers to leverage hardware acceleration and interact with external resources that are not constrained by the borrow checker. These features can also improve performance through bypassing certain runtime assertions, like bounds checks.
@@
However, if unsafe code is used inappropriately, it can reintroduce the same types of safety issues that Rust was designed to prevent, such as data races and accesses-out-of-bounds. Rust’s aliasing restrictions also introduce new forms of undefined behavior. Though references are constrained by the borrow checker, they can also be cast into raw pointers and used freely in unsafe contexts. Rust’s compiler is allowed to assume that programs which use raw pointers still adhere to the rules of Rust’s aliasing model. If a program breaks these rules, then it may be optimized incorrectly, introducing differences in behavior that can constitute security vulnerabilities (Crichton and Vranken, 2023).
2.1. Best Practices & Tooling
Several best practices can help developers mitigate the risks of using unsafe code. When implementing an unsafe feature, developers are encouraged to follow the “interior unsafety” pattern by keeping it minimal, well documented, and hidden beneath a safe interface (Qin et al., 2020). Ideally, this interface should make it impossible for users to trigger undefined behavior from safe code. However certain safety properties are difficult or impossible to encode into Rust’s type system without leveraging external verification tools. If a function or trait places the burden of correctness on its users, then it should be labelled unsafe and documented with any necessary invariants. These and other guidelines are provided by multiple community resources. The Rust Reference (Rust Community, 2023a) and The Rust Programming Language (Klabnik and Nichols, 2022) each include relevant sections on unsafe code, but the Rustonomicon (Rust Community, 2023b) provides the most comprehensive overview. The Unsafe Code Guidelines Reference (Rust Community, 2023c) was intended to be an in-depth guide to the semantics of unsafe features, but it is incomplete and mostly unmaintained. However, the “Issues” section of its repository is still actively used to discuss Rust’s semantics.
In addition to these best practices, dynamic analysis is also regularly used for validating unsafe code. Miri (Rust Community, 2024f), a Rust interpreter, is currently the only method for detecting violations of Rust’s evolving aliasing models, so it is seen as the “de facto” (Gäher et al., 2024) bug finding tool for Rust. Currently, Miri can detect violations of two different models of Rust’s semantics. Stacked Borrows (Jung et al., 2019) was the first iteration of Rust’s aliasing rules, and it is enabled by default. Developers can also switch to using the newer Tree Borrows (Villani et al., 2023) model by providing a configuration flag. Neither of the two models is canonical. Shuttle (Rust Community, 2024h) and Loom (Rust Community, 2024e) are Rust-specific fuzzing tools that permute the scheduling of threads to help detect concurrency bugs. Rust is also compatible with several language agnostic tools, including libFuzzer (LLVM Project, 2022), the LLVM project’s sanitizers (LLVM Project, 2024), and Valgrind (Nethercote and Seward, 2007).
Static analysis is also useful in this context. Clippy (Rust Community, 2024d), a linting tool, can hint at improper use of unsafe features, but it is not a true bug finding tool. The Rust ecosystem does not yet have a de facto solution for static bug finding, but several research efforts have produced promising results. Rudra (Bae et al., 2021) can detect use-after-free errors when a Rust application panics, while MIRChecker (Li et al., 2021) can detect panics and duplicated ownership. The Rust Analysis Platform (RAP) (Cui et al., 2023a, 2024) can detect memory leaks and use-after-free errors. MIRAI aims to provide a unified solution for both static analysis and verification with abstract interpretation (Endor Labs, Rust Community, 2024). Other extensions leverage community knowledge to help enforce safety properties. These have typically been developed as plugins for cargo, Rust’s build tool. Both cargo-audit (Rust Community, 2024g) and cargo-deny (Rust Community, 2024a) detect security vulnerabilities in a project’s dependencies; cargo-deny is more extensible. The cargo-geiger (Rust Community, 2024b) plugin provides statistics on unsafe code, and cargo-vet (Rust Community, 2024c) can be used to track if code has been reviewed by a trusted party.
Verification has been a uniquely active area of research within the Rust community. Here, we provide a high-level overview using several representative tools. Each category of tooling has a different level of compatibility with unsafe code. The first category includes Prusti (Astrauskas et al., 2019), Creusot (Denis et al., 2022), and Aeneas (Ho and Protzenko, 2022), each deductive verifiers, and Flux (Lehmann et al., 2023), a refinement type checker. These tools rely on Rust’s aliasing invariants to encode the properties of programs into their logic, so they do not support unsafe code by construction. The second category includes RefinedRust (Gäher et al., 2024) and Gillian-Rust, which exclusively target unsafe code. RefinedRust encodes encodes programs into a separation logic defined in Iris, which is automatically verified using the Lithium engine. Gillian-Rust relies on symbolic execution. Both approaches are informed by the RustBelt (Jung et al., 2017) model. Remaining tools are targeted at verifying arbitrary Rust programs, but they use unique approaches that provide limited support for unsafe features. Verus (Lattuada et al., 2023), a deductive verifier, uses linear ghost types to verify heap accesses through raw pointers. Kani, a model checker (VanHattum et al., 2022), is capable of detecting spatial and temporal memory safety issues. Though each of these tools have varying levels of support for unsafe code, none can validate programs against Miri’s evolving definitions of Rust’s aliasing model, and only Kani has support for foreign function calls.
3. Related Work
Table 1. A comparison table outlining how we contribute to prior studies of unsafe Rust. The column “Focus” indicates whether a work exclusively focused on unsafe code (direct) or studied the Rust ecosystem more broadly (indirect).
{tblr}
colspec=Q[l,m]Q[l,m]Q[l,m]$Q[c,m]Q[c,m]$ Work & Focus Methodology RQ1 RQ2 RQ3 RQ4
Astrauskas (Astrauskas et al., 2020) Direct Code Analysis
Cui (Cui et al., 2023b) Direct Code Analysis, Survey (n=50)
Evans (Evans et al., 2020) Direct Code Analysis, Survey (n=20)
Fulton (Fulton et al., 2021) Indirect Interviews (n=16), Survey (n=178)
Höltervennhoff (Höltervennhoff et al., 2023) Direct Interviews (n=26)
McCormack (McCormack et al., 2024) Direct Code Analysis
Ozdemir (Ozdemir, 2016) Direct Code Analysis
Qin (Qin et al., 2020) Direct Code Analysis
Xu (Xu et al., 2021) Direct Code Analysis
Zeng (Zeng and Crichton, 2018) Indirect Content Analysis, Interviews (n=3)
Zhu (Zhu et al., 2022) Indirect Content Analysis, Survey (n=101)
Zheng (Zheng et al., 2023) Direct Code Analysis
[…]
[c=1]r Combined:
{tblr}colspec=Q[c,m]Q[l,m] Label Definition
Provides relevant evidence, but does not directly answer our research question.
Gives a partial answer to our research question, but we provide additional parts.
Answers the same research question—we evaluate if their results generalize to a larger population.
Answers the same research question.
Does not address our research question.
Prior studies have used a variety of methods to investigate how Rust developers use unsafe code. Here, we provide an overview of results that are relevant for each of our research questions. In Section 3.1, we describe results from studies that examined how unsafe code is written. We describe studies that engaged directly with Rust developers in Section 0. In Section 3.3, we classify every relevant prior study and outline how we make novel contributions in our answers to each research question.
3.1. Studies of Code
Distribution
Surveys indicate that unsafe code is prevalent throughout the Rust ecosystem, but developers tend to use it minimally and hide it beneath safe interfaces. In Rust, packages and libraries are referred to as “crates,” and they are published on crates.io. In July of 2016, Ozdemir (Ozdemir, 2016) found that 29% of all published crates used unsafe features, but most unsafe blocks were relatively small and tightly scoped to include only unsafe operations. Evans et al. (Evans et al., 2020) observed similar patterns in September of 2018. The number of published crates had grown by more than ten times, but 29% still directly used unsafe code. They also observed that 38% of crates did not use any unsafe features, but they directly or indirectly depended on third-party crates that did. In January of 2020, Astrauskas et al (Astrauskas et al., 2020) found that the proportion of crates using unsafe features had shrunk to 23.6%. However, most instance of unsafe code were relatively straightforward and encapsulated. The majority of unsafe functions called by libraries were standard (e.g. not closures, function pointers, or trait functions) and defined within the same crate or provided by Rust’s standard library. More than 80% of unsafe functions were publicly visible, but most were exposed by crates that provided raw bindings to foreign libraries.
Use Cases
Studies have also examined applications to identify use cases for unsafe code. Qin et al. (Qin et al., 2020) studied a random sample of 600 instances of unsafe code from 10 popular Rust libraries, as well as 250 instances within safe encapsulations provide by Rust’s standard library. They identified three use cases for these operations; 42% were related to interoperation, 22% were for improving performance, and 14% involved sharing memory between threads. Astrauskas et al. (Astrauskas et al., 2020) used “anecdotal examples” and snippets of source code to create a taxonomy of six purposes for using unsafe code. These included implementing data structures with complex sharing, overcoming the incompleteness of Rust’s type system, emphasizing contracts and invariants, calling foreign functions, using intrinsics or inline assembly, and improving performance.
Bugs & Vulnerabilities
Discrepancies between the signatures of safe APIs and their interior unsafe implementations have been a significant source of bugs and vulnerabilities within the Rust ecosystem. Qin et al. (Qin et al., 2020) examined 170 bugs and vulnerabilities from ten high-profile Rust libraries. The Memory safety errors that they encountered were most frequently triggered from within safe contexts due to errors in unsafe implementations elsewhere. They also identified 19 examples of improper encapsulations of unsafe code in both the third-party libraries that they sampled and within Rust’s standard library. Xu et al. (Xu et al., 2021) exclusively studied registered vulnerabilities and identified similar errors in the encapsulations of unsafe code. Zheng et al. (Zheng et al., 2023) conducted the broadest empirical study of security issues in the Rust ecosystem to date. They examined 433 unique vulnerabilities that were discovered between 2014 and 2022, identifying 300 affected repositories. The portions of code that were affected by a vulnerability were significantly more likely to include unsafe blocks or functions. The overwhelming majority of all vulnerabilities affected safe functions.
It can be difficult to correctly implement an encapsulation because unsafe implementations require a variety of safety invariants. Cui et al. (Cui et al., 2023b) manually reviewed 416 unsafe APIs in Rust’s standard library and identify 19 categories of preconditions and postconditions. They evaluated their taxonomy with a survey of 90 Rust developers, and also categorized 186 vulnerabilities based on the safety properties that were violated. The most common violations were related to incorrect uses of Send or Sync traits, incomplete initialization, or accesses-out-of-bounds.
Interoperation
Interoperation has consistently been one of the most prevalent use cases for unsafe code. Astrauskas et al. (Astrauskas et al., 2020) found that 44.6% of all publicly-exported unsafe functions were static bindings to foreign functions. Other surveys of the Rust ecosystem have similar conclusions; Ozdemir (Evans et al., 2020) found that 30% of unsafe blocks were entirely dedicated to foreign function calls, and Evans et al. (Evans et al., 2020) found that 22.5% of unsafe function calls were to foreign functions. Few studies have examined errors that occur due to the interaction between Rust and other languages. Cui et al. (Cui et al., 2023b) explicitly excluded foreign functions from their investigation of safety properties. Xu et al. (Xu et al., 2021) found 12 examples of bugs and vulnerabilities that involved foreign function calls. They were related to application-specific invariants, data races, invalid alignment, and inconsistent layouts. The authors provided few details on the nature of these bugs, indicating that they were “straightforward” and not Rust-specific.
Only two studies have identified bugs that meaningfully involve Rust’s unique aliasing restrictions. In 2023, a set of anonymous authors categorized 22 bugs from five popular Rust encapsulations for foreign libraries (Anonymous, 2023). Most of the errors that they investigated involved incorrect handling of exceptions and heap objects that were improperly shared across foreign boundaries. They identified two bugs related to Rust’s aliasing restrictions, but they did not describe these bugs in terms of Rust’s operational semantics. McCormack et al. (McCormack et al., 2024) are the only group to have studied violations of Rust’s aliasing models in multi-language applications. They created a prototype extension to Miri (Rust Community, 2024f) and used it to find 48 bugs in Rust libraries that called foreign functions. Half of these bugs were violations of Rust’s Tree Borrows (Villani et al., 2023) aliasing model. A few were found in high-profile libraries, including one maintained by the Rust project.
3.2. Studies of Developers
Content Analysis
Studies of online discourse have also provided relevant insights into encapsulation and interoperation. Zeng and Crichton (Zeng and Crichton, 2018) studied the challenges associated with adopting Rust. They interviewed three Rust contributors and analyzed a set of articles and comments posted on Hacker News and the Rust subreddit (/r/rust). They indicated that it was more common for developers to discuss rewriting entire applications in Rust instead of porting them incrementally, suggesting that the difficulties of interoperation may prevent developers from taking this route. They also identified that developers who struggled with Rust’s aliasing rules were often pointed toward safe encapsulations of unsafe operations. Zhu et al. (Zhu et al., 2022) observed a similar phenomenon on Stack Overflow. They used a mixed-methods approach to understand how developers reason about Rust’s aliasing restrictions. In the first stage of their investigation, they examined a random sample of 100 Stack Overflow questions to identify situations where developers asked for help fixing code that had been rejected by the borrow checker. Only three of the 110 borrowing errors in these questions were fixed by writing unsafe code. Developers were referred to safe alternatives in every other situation.
Qualitative & Mixed-Methods Studies
Fulton et al. (Fulton et al., 2021) studied the benefits and drawbacks of adopting Rust as a general-purpose systems programming language. They used an exploratory, mixed-methods design that is similar to ours. First, they interviewed 16 developers in industry who had advocated for using Rust. Then, they distributed a community survey covering the same topics as the interviews, and it received 178 responses. The majority of their participants struggled with Rust’s steep learning curve, but error messages from the compiler and community resources made it easy for them to find solutions to their problems. Participants felt somewhat more productive using Rust, and they were more confident that their implementations were correct. Fulton et al. (Fulton et al., 2021) only briefly discussed unsafe features. Their participants typically used unsafe code to call foreign functions, improve performance, and interact directly with hardware or the operating system.
Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only other group that we are aware of who have used qualitative methods to study how developers use unsafe code. They interviewed 26 developers with experience using unsafe Rust. Except for three who piloted the interview protocol, all of their participants had contributed unsafe code to open-source projects included in the unofficial “Awesome Rust” list (Community, 2022) of applications. We cover similar topics, but our research questions reflect different, complementary goals. They focused on understanding how developers contend with the security implications of using unsafe code, while we are focused on interoperation and tool use. Most of their participants relied on common sense and test coverage to evaluate the security of their code, and few had contributed to projects with formal policies on security or code review. However, most indicated that that they would review unsafe code more carefully than safe code, ensuring that it upheld all necessary safety invariants. Few participants had ever encountered security issues caused by unsafe code, and none of the issues that participants did experience were reported as critical vulnerabilities. We provide additional details on the results that are relevant for answering our research question in Section 0 and throughout Section 0.
3.3. Our Contributions
In Table 0, we use multiple descriptors to classify each of the relevant studies that we presented in previous sections. We label the methods used by each study as code analysis, content analysis, surveys, and interviews. We define code analysis to include mining software repositories or gathering information from bug reports and vulnerabilities. Content analysis includes examining any other type of preexisting online content, such as forum posts, blog posts, articles, or online discussions, to gather insights about Rust developers. Studies with a direct focus on unsafe code are similar to ours in that each of their core research questions are targeted at understanding how developers use unsafe features. Studies with an indirect focus are more broadly concerned with Rust but also happen to provide relevant insights for answering our research questions.
We also use several labels to indicate whether each study can fully answer our research questions. The empty circle ‘ ’ indicates that a study does not answer a question. The label ‘ ’ indicates that a study provides relevant evidence, but it does not address the same question. The label ‘ ’ indicates that a study partially answers our question, but we provide missing parts. The label ‘ ’ indicates that a study fully answers our question, but it does not provide findings that generalize to a population that is comparable to our survey population. Now, we describe our rationale for the labels that we chose and we indicate how we provide new results in our answers to each question.
Interoperation (RQ1)
Studies of bugs, vulnerabilities, and the distribution of unsafe code all indicate that foreign function calls are a potentially significant unchecked source of undefined behavior in the Rust ecosystem. However, none of these studies provide insights into the challenges that developers face when implementing these applications. Most of the participants studied by Fulton et al. (Fulton et al., 2021) indicated that C and Python were easy to interoperate with, but C++ was difficult to handle due to complex aliasing. Participants interviewed by Höltervennhoff et al. (Höltervennhoff et al., 2023) also had mixed experiences with interoperation; four described it as being easy, but two struggled with “memory management, e.g. between programming languages.” These studies can only partially answer our research question ( ), since neither provides a detailed description of the patterns that lead to these difficulties, other than that they involved aliasing and initialization (Höltervennhoff et al., 2023; Fulton et al., 2021). We have an explicit focus on interoperation as one of our core research questions, and we provide a missing perspective on the challenges of using Rust in multi-language applications.
Tooling (RQ2)
Rust is a relatively new programming language, so the majority of research efforts have focused on developing new tools instead of evaluating existing ones. Zeng and Crichton (Zeng and Crichton, 2018) hypothesized that a lack of promotion for tooling could prevent Rust from being widely adopted as a systems language. However, their work was published prior to the creation of most Rust-specific bug finding tools. Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only other group that we are aware of who have examined how developers use tooling to assist with writing unsafe code. Key development tools were somewhat underused among their interview population—only 7 of their 26 participants had mentioned using Miri. Since this is a small sample size, it is unclear whether these results are indicative of a broader trend in the Rust community, so they can only partially answer our second research question ( ). We also cover debuggers, static verifiers, and tools for generating bindings to foreign functions, which were not mentioned by Höltervennhoff et al. (Höltervennhoff et al., 2023).
Motivations (RQ3)
Several prior studies have conflated developers’ motivations for using unsafe code with features-oriented use cases. For example, Astrauskas et al. (Astrauskas et al., 2020) include both performance and compiler intrinsics as “motivations” for using unsafe code. In their survey of 20 developers, Evans et al. (Evans et al., 2020) defined performance, ease-of-use, and a variety of specific operations as “reasons” for using unsafe code. Fulton et al. (Fulton et al., 2021) include both performance, interaction with hardware, and foreign function calls as “uses” for unsafe code. We define “motivations” as feature agnostic reasons for using unsafe code. This perspective is similar to Höltervennhoff et al. (Höltervennhoff et al., 2023), who identified necessity and performance, and efficiency as “motivations,” but the separation between features and higher-level reasons was not an explicit design consideration of their study. When viewed together, these results answer our third research question ( ). However, we provide a novel contribution by replicating their findings and evaluating whether they generalize to a large population of developers. Additionally, Höltervennhoff et al. (Höltervennhoff et al., 2023) study developers’ motivations as part of examining the broader topic of the “common practices for deciding for and implementing” unsafe code. We dedicate an entire research question to this topic, so we are able to provide significantly more detail on how specific design decisions relate to each motivation.
Encapsulation (RQ4)
Surveys of the Rust ecosystem suggest that developers typically follow best practices for minimizing and encapsulating unsafe code, but they struggle with ensuring that their interfaces are correct in all situations. Höltervennhoff et al. (Höltervennhoff et al., 2023) are the only prior study to have examined developers experiences with designing applications that include unsafe code, and their results reflect the patterns observed surveys of code and bug reports. The majority of their participants would avoid using unsafe code whenever possible. When unsafe code was necessary, they ensured that it was encapsulated beneath a safe interface of their own design or provided by an external library. However, the majority of participants felt that it was difficult to write unsafe code correctly; they had a “nagging feeling about it”, and half were unsure or undecided about its correctness. Similar to RQ3, these results also answer our research question ( ), but we provide a unique contribution by evaluating whether their results generalize to our survey population. We also use the set of motivations we defined in our answer to RQ3 to investigate developers’ motivations for exposing unsafe APIs, which is not addressed by Holtervennhoff et al. (Höltervennhoff et al., 2023).

### v4

# A Mixed Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

@@ Abstract.
The Rust programming language restricts aliasing to provide static safety guarantees. However, in certain situations, developers need to bypass these guarantees by using a set of “unsafe” features. When these features are used incorrectly, they can reintroduce the types of safety issues that Rust was designed to prevent. We seek to understand how current development tools can be improved to better assist developers who find it necessary to interact with unsafe code. To that end, we study how developers reason about foreign function calls, the limitations of the tools that they currently use, their motivations for using unsafe code, and how they reason about encapsulating it. We conducted a mixed methods investigation consisting of semi-structured interviews with 19 developers, followed by a survey that reached an additional 160 developers. Our participants were motivated to use unsafe code when they perceived that there was no alternative, and most claimed that they would avoid using it. However, limited tooling support for foreign function calls made participants uncertain about their design choices, and certain foreign aliasing and concurrency patterns were difficult to encapsulate. To overcome these challenges, developers will need analysis tools that can find Rust-specific forms of undefined behavior within multilanguage applications.
$c$s: Software and its engineering
@@ 1. Introduction
The Rust programming language has made rapid headway as a safe alternative for systems programming due to its static safety guarantees. However, it is still the case that critical systems are written in languages that do not provide inherent safety guarantees (Rust for Linux, 2026; Jansens, 2023). Rust developers who interoperate with these applications must use a special subset of “unsafe” features, including calling foreign functions and accessing memory through raw “C-style” pointers. These features can only be used within a block or function marked with the unsafe keyword.
@@
To minimize the risks of unsafe code, the Rust community advocates for keeping it minimal, well documented, and hidden beneath a safe interface (Astrauskas et al., 2020). However, these practices alone are not enough to ensure that Rust delivers on its promise of static safety. Errors in safe encapsulations of unsafe implementations have been a significant source of bugs and security vulnerabilities in the Rust ecosystem (Qin et al., 2020; Xu et al., 2021; Zheng et al., 2023). Applications that rely on interoperation or direct access to system resources may expose developers to sources of unsafety that cannot easily be minimized. Making matters worse, Rust’s static and dynamic semantics are still evolving (Villani et al., 2025; Rémy Rakic, 2023), so few tools are capable of fully testing or verifying these applications.
@@
Interoperation is a common use case for unsafe code (Evans et al., 2020; Astrauskas et al., 2020; Fulton et al., 2021; Höltervennhoff et al., 2023), but most Rust-specific development tools lack support for foreign function calls. Safe interoperation will be crucial for Rust’s future, since both industry and regulatory stakeholders have advocated for adopting Rust as a replacement for existing systems languages (Wallach, 2024; Rebert and Kern, 2024). Automated translation tools are increasingly effective (Larsen, 2018), but for the moment, interoperation is the most practical method for adopting Rust within existing, large-scale systems (Rebert and Kern, 2024). It can be challenging to use Rust correctly in this context, because it has significant differences from other languages (Höltervennhoff et al., 2023; McCormack et al., 2025b). A detailed account of these challenges is missing, and we need this to be able to develop effective tooling for safe interoperation. To provide this perspective, we ask the following research questions:
RQ1 (Interoperation).
@@
How do Rust developers reason about memory safety across foreign function boundaries?
RQ2 (Tooling).
What tools do Rust developers use when contributing to applications that include unsafe code, and how could tooling be improved?
@@
Developers who interoperate with Rust will need to use a variety of unsafe features in addition to foreign function calls to be able to reconcile the differences between Rust and other languages. Having a broader understanding of what motivates Rust developers to use unsafe code in any capacity and how they reason about soundness will be helpful to identify if there are factors beyond these differences that impact the safety of multilanguage applications. However, most large-scale studies on these topics have analyzed patterns in source code instead of directly engaging with developers (Evans et al., 2020; Astrauskas et al., 2020). Interviews and surveys have provided a missing qualitative perspective, but tend to have a limited focus on unsafe code (Fulton et al., 2021), a small sample size (Evans et al., 2020), or different goals (Höltervennhoff et al., 2023). In addition, to serve as guidance for the design of future tools, we seek to generalize prior qualitative results to a broader population of developers. To that end, we ask the following additional research questions:
RQ3 (Motivations).
What are Rust developers’ motivations for using unsafe code?
RQ4 (Encapsulation).
How do Rust developers reason about encapsulating unsafe code?
@@
We provide the results of an exploratory, mixed method study that addresses the research questions above. We conducted semi-structured interviews with 19 developers and used the results to create a community survey that had 160 valid responses. Most of our participants used unsafe code to call foreign functions, and they described several aliasing and concurrency patterns that made interoperation difficult. Many developers had used Miri (Jung et al., 2026), a Rust interpreter, to find bugs, but most were deterred from using it again due to its slow performance and lack of support for key features. Participants prioritized keeping unsafe code minimal, encapsulated, and well documented, but rarely audited their dependencies and relied on ad hoc reasoning to justify their decisions. Expanded tooling support for multilanguage applications and improved documentation will be necessary to support developers in maintaining Rust’s safety guarantees across foreign function boundaries.
@@
In Section 0, we describe the role of unsafe code in the Rust ecosystem and the challenges associated with interoperating with other languages. We describe our methodology in Section 0, and we indicate threats to validity in Section 0. We present our results in Section 0 and compare our findings with prior work in Section 0. We discuss the implications of our results in Section 0 before concluding in Section 0. Our replication package is available via Zenodo (McCormack et al., 2026).
@@ 2. Background
Rust’s safety guarantees are based on the notion of ownership. Values with copy semantics do not have an owner and can be copied freely, while values with move semantics have a unique owner. Assignment transfers ownership from one alias to another, but ownership can also be borrowed by creating a reference. Each reference has a lifetime, which is the duration of the program where it is used. Rust’s borrow checker enforces that a value can have either a unique mutable reference or many immutable references, but not both simultaneously (Crichton, 2020). A value’s type can implement traits that define how it behaves. Certain traits affect a value’s capability to be owned and borrowed. Values with the Copy trait have copy semantics, values that are Send can be moved between threads, and values that are Sync can be borrowed by multiple threads.
@@
Rust’s aliasing rules are conservative and are not universally compatible with certain idioms of systems programming. When these restrictions become burdensome, developers can use the unsafe keyword to enable features that are not constrained by the borrow checker. Within an unsafe context, users can dereference raw pointers, execute inline assembly instructions, modify static mutable state, implement unsafe traits, and call unsafe functions (Klabnik et al., 2026). There are a variety of uses for these features (Astrauskas et al., 2020). In certain situations, unsafe code can be necessary to implement design patterns that are fundamentally incompatible with the restrictions of safe references, such as shared mutable state. Foreign function calls, inline assembly, and intrinsics allow developers to leverage hardware acceleration and interact with external resources. These features may also improve performance by skipping certain run-time assertions, such as bounds checks.
@@
However, if unsafe code is used inappropriately, it can reintroduce the same types of safety issues that Rust was designed to prevent, such as data races and accesses out-of-bounds. Rust’s aliasing restrictions also introduce new forms of undefined behavior. Safe references are constrained by the borrow checker, but they can also be cast as raw pointers, which can only be used in unsafe contexts. The Rust compiler is allowed to assume that programs which use these features are still following Rust’s aliasing rules. Programs that break these rules may be optimized incorrectly, introducing differences in behavior that can constitute security vulnerabilities.
Several best practices can help mitigate the risks of using unsafe code. When implementing an unsafe feature, developers are encouraged to follow the “interior unsafety” pattern by keeping it minimal, well documented, and hidden beneath a safe interface (Qin et al., 2020). Ideally, this interface should make it impossible for users to trigger undefined behavior from safe code. However, certain safety properties are difficult or impossible to encode into Rust’s type system without leveraging external verification tools. Miri (Jung et al., 2026), a Rust interpreter, is currently the only tool that can detect when unsafe operations lead to violations of Rust’s newest aliasing rules, so it is typically seen as the “de facto” (Gäher et al., 2024) bug-finding tool for applications that use unsafe code. Currently, Miri can detect violations of two different models of Rust’s semantics. The Stacked Borrows (Jung et al., 2019) model was the first iteration of these rules; it is enabled by default. Developers can also switch to using the newer Tree Borrows (Villani et al., 2025) model by providing a configuration flag. Neither of the two models is canonical.
These practices are standard for the Rust community, but Rust is also increasingly being adopted in interoperation with preexisting large-scale C and C++ applications, such as Chromium (Jansens, 2023) and the Linux Kernel (Rust for Linux, 2026). These applications will expose Rust components that may be “safe” in isolation to significant external sources of unsafety, which will not be as easy to minimize or document. In addition to these challenges, Miri does not fully support finding undefined behavior in this context. Miri can execute and trace certain foreign functions from shared libraries, but it cannot detect aliasing violations that are triggered in foreign code, and its slow performance prevents it from being used at scale.
These obstacles will make it difficult for developers to avoid introducing undefined behavior in multilanguage Rust applications. Projects such as the DARPA’s TRACTOR program (Wallach, 2024) and the Rust/C++ Interoperability Initiative (Bauman and The Rust Foundation, 2024) seek to provide new development tools to assist with these challenges. We need a broader understanding of how Rust developers use unsafe code to be able to design interventions that are effective in practice, since developers who use Rust in this context will necessarily engage with a variety of unsafe operations in order to reconcile differences in semantics across language boundaries.